### v1

# Beyond Community Notes: A Framework for Understanding and Building Crowdsourced Context Systems

Abstract.
Social media platforms are increasingly developing features that display crowdsourced context alongside posts, modeled after X’s Community Notes. These systems, which we term Crowdsourced Context Systems (CCS), have the potential to reshape our information ecosystem as major platforms embrace them as alternatives to top-down fact-checking. To deeply understand the features and implications of such systems, we perform a systematic literature review of existing CCS research and analyze several real-world CSS implementations. Based on our analysis, we develop a framework with three distinct components. First, we present a theoretical model to help conceptualize and define CCS. Second, we identify a design space encompassing six key aspects of CCS: participation, inputs, curation, presentation, platform treatment, and transparency. Third, we identify key normative implications of different CCS design and implementation choices. Our framework integrates these theoretical, design, and ethical perspectives to establish a foundation for future human-centered research on Crowdsourced Context Systems.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
Refer to caption Refer to caption :
Figure 1. Screenshots of a Community Note displayed on Instagram (left) and X (right).
1. Introduction
Social media platforms play an important role as sites of entertainment (Herder and Staring, 2024), news consumption (Chouaki et al., 2024), and interpersonal connection (Preece, 1999). At the same time, these platforms must contend with problems stemming from harmful user-uploaded content, which can take the form of abuse, spam, or misinformation (Scheuerman et al., 2021). Specifically, platforms have tried various forms of content moderation (Grimmelmann, 2015; Lampe and Resnick, 2004; Jhaver et al., 2019) to deal with misinformation, including centralized fact-checking efforts to label or remove it (Gillespie, 2018), as well as distributed approaches where users are empowered to tackle these challenges themselves (Seering, 2020). While these approaches have had some success, they have also evoked backlash and accusations of bias from both users and lawmakers, particularly on the political right (Jiang et al., 2020; Johnson, 2023). Recently, a new approach has emerged: beginning with X/Twitter, major social media platforms are increasingly introducing systems that crowdsource “helpful context” to display alongside potentially problematic posts (Coleman, 2021) (e.g. Figure 1). These systems, which we term Crowdsourced Context Systems (CCS), have seen growing adoption and impact, with some of the world’s largest platforms, including Meta (Facebook, Instagram, and Threads), TikTok, and YouTube all recently launching their own implementations. Meta has gone one step further, announcing that their CCS will replace their fact-checking program in the U.S. (Kaplan, 2025).
As CCS adoption grows, these systems have the potential to reshape our information ecosystem. Early studies into such systems have produced promising results, finding that CCS can increase user informedness and decrease the spread of misinformation (Wojcik et al., 2022; Chuai et al., 2024b; Drolsbach and Pröllochs, 2023), although it is still unclear whether CCS can replace centralized fact-checking, as some platforms have done or may seek to do (Kaplan, 2025). As CCS develop, novel implementation questions also come to the fore, including how to design these systems most effectively, for example, to provide reliable context quickly enough to prevent the spread of the most viral content (Chuai et al., 2024b, a; Renault et al., 2024). At this critical moment of change in the information ecosystem, it is important to understand how CCS function, how we should evaluate them, and how these systems may be better designed.
This paper makes two contributions towards the goal of better understanding CCS. Our first contribution is a systematic literature review of existing research studying implementations of CCS. Our review finds significant research into CCS contributor behavior and the impact of displaying notes. However, the review highlights the fact that certain aspects of these systems are underexplored by existing research. We thus set out to identify additional facets of these systems that may affect their impact. Our second contribution is an analysis and comparison of real-world implementations of CCS by several major social media platforms, in order to create a framework to guide future HCI research on these novel and important systems. Our framework is composed of three parts. First, we offer a theoretical model of a CCS that identifies these systems’ core characteristics and what sets them apart from prior crowdsourced systems. Our model posits that while CCS use familiar moderation tools like crowdsourced information and context labels, they are unique in how they are natively integrated into platforms with the explicit goal of fostering user informedness. Second, we outline a design space (MacLean et al., 1996) for CCS in order to taxonomize the important design decisions made in their implementations. Design spaces are a conceptual tool used in HCI research that can guide work on complex socioetechnical systems by structuring the space of design possibilities (Zhang et al., 2024; Lee et al., 2024; Card et al., 1990). We group these design decisions into six high-level aspects of CCS that combine and interact to shape their impact: participation, inputs, curation, platform treatment, and transparency. Third, we identify key normative implications of these systems in order to make sense of the ethical stakes involved in their design. As systems with the power to shape the information ecosystem, we emphasize three interconnected concerns: user informedness, distribution of power, and fairness. We discuss the role of CCS within the broader conversation about content moderation and establish a foundation for future human-centered research on Crowdsourced Context Systems.
2. Background
In January 2021, Twitter (now X) launched a U.S. pilot for the first CCS, which they called Birdwatch, presenting it as, “a community-based approach to misinformation” (Coleman, 2021). This system allowed a subset of users to suggest notes to, “provide informative context” about Tweets (posts) on the platform. The same subset of users then rated these suggested notes and these ratings were used as the inputs to a note curation algorithm that aimed to select notes that would be helpful to a broad audience. In this pilot, these notes were only visible on a separate site, but Twitter announced their plan, “to make notes visible directly on Tweets for the global Twitter audience, when there is consensus from a broad and diverse set of contributors.”
In March 2022, Twitter expanded the pilot, announcing that now, “a small (and randomized) group of people on Twitter in the US will see Birdwatch notes directly on some Tweets” (Coleman, 2022). Twitter was purchased by Elon Musk in October 2022, renamed X, and soon Birdwatch was renamed Community Notes. In December 2022, the pilot phase ended and users all over the world began seeing selected notes directly alongside Tweets (Community Notes [@CommunityNotes], 2022). Since then, the system has continued to evolve. A key feature of X’s Community Notes is its transparency: the note curation code is open source [^1] https://github.com/twitter/communitynotes/tree/main, as are the crowdsourced notes and ratings data collected by the system [^2] https://x.com/i/communitynotes/download-data.
In June 2024, YouTube announced a pilot for a similar system, which they called Notes (Team, 2024). In January 2025, Meta announced a pilot for their own implementation, also called Community Notes, which would launch on Facebook, Instagram, and Threads (Kaplan, 2025). Finally, TikTok announced their own pilot in April 2025, which they called Footnotes (Presser, [n. d.]). According to the platforms, all of these new systems are based on X’s open source implementation of the note curation algorithm. Unlike X’s implementation, though, none of the new systems are open source or open data. In fact, while all of these systems serve similar purposes, they differ from each other in crucial ways due to unique design decisions made by each platform. Since these systems are increasingly prominent features of some of the most important social media platforms, it is critical to define, understand, and reason about them so the HCI community can better inform their implementation and evaluate their impact.
3. Related Work
While CCS are a relatively new innovation, these systems relate to and touch on multiple bodies of work from the HCI literature, including: research on discourse and deliberation systems, research on fact-checking infrastructures, and research on crowdsourced responses to misinformation. We describe these areas here; later in the paper (Section 6) we provide a detailed review of the literature that is specifically addressing CCS.
In a democracy that values free speech, public participation in decision-making via deliberation is considered a social good (Small, 2021). One classic concern of HCI research has thus been how to design systems that can support public deliberation among users with diverse, possibly conflicting, perspectives (Kriplean et al., 2012). One such approach was recently termed bridging systems, or systems that, “increase mutual understanding and trust across divides, creating space for productive conflict, deliberation, or cooperation” (Ovadya and Thorburn, 2023). Such systems can be built with various goals in mind, such as finding common ground between peace-builders in armed conflict (Konya et al., 2025), or facilitating participatory approaches to allocate public funding (Yang and Bachmann, 2025). CCS can be viewed as highly distributed deliberation systems that aim to allow more informed public participation in social media. Community Notes, X’s CCS implementation, is one such system. Indeed, according to its creators, Community Notes was designed to identify crowdsourced content (notes), “that both inform understanding (decrease propensity to agree with a potentially misleading claim) and are seen as helpful by a diverse population of users” (Wojcik et al., 2022). CCS thus reflect a long-standing HCI concern with designing systems that support civic deliberation, adapting these ideas to meet the scale, speed, and additional challenges of major social media platforms.
CCS also align with another classic goal of HCI research, which is to help users assesses the veracity of information that they encounter on social media. Emerging from journalistic practice, one approach to address this issue is to rely on external fact-checking (Walter et al., 2020; Young et al., 2018), where independent organizations verify public claims and publish their determinations. While these techniques can have positive effects, they are time-consuming and traditionally require a large amount of professional labor (Hassan et al., 2015; Juneja and Mitra, 2022), and thus may struggle keep up with the speed at which content is published on the internet (Micallef et al., 2022). As a result, researchers have explored more scalable approaches, such as automated tools to support human fact-checkers (Warren et al., 2025), as well as fully-automated approaches (Guo et al., 2022; Hassan et al., [n. d.]). While external fact-checks are often found on websites different from where the claim originated, such as Snopes [^3] https://www.snopes.com/ or Politifact [^4] https://www.politifact.com/, social media platforms may also provide their own fact-checks, which can be presented directly alongside user-posted content (How, [n. d.]). Studies have investigated both the most effective ways to present these fact-checks in order to shift belief (Clayton et al., 2020) and also the psychological characteristics that make some users reluctant to engage with them (Tanaka et al., 2023). In addition to issues of efficacy, and the large cost of running such fact-check programs, social media platforms have struggled with allegations and perceptions that their fact-checks are biased against certain political ideologies (Cotter et al., 2022).
While fact-check systems traditionally rely on a central authority, alternative approaches to dealing with information of questionable quality make use of the distributed expertise of a crowd (Pennycook and Rand, 2019; He et al., 2025). Crowd-based fact-checking is one such approach, which despite being cheaper and faster than professional fact-checking, has often been found to be similarly accurate and effective (Allen et al., 2021; Martel et al., 2024; Zhao and Naaman, 2023). HCI research has explored the practices of those who participate in crowdbased fact-checking (He and He, 2022), their potential biases (Draws et al., 2022), and how fact-checking tasks may be best distributed to a crowd (Bhuiyan et al., 2020). Beyond fact-checking, there are other crowd-based approaches to handling misinformation, such as digital juries (Fan and Zhang, 2020; Zhao and Hobbs, 2025), distributed moderation through up/down votes (Lampe and Resnick, 2004; Epstein et al., 2020), and community self-governance (Seering, 2020). CCS share much in common with these previous crowdsourced fact-checking systems, but they have a novel set of processes and structures that make CCS stand out as a new type of interactive systems. Informed by these past bodies of research, we next define what makes these new systems unique.
Figure 2. Model of a Crowdsourced Context System. Users create posts, context contributors create notes about those posts, and ratings contributors create ratings for those notes. Based on these ratings, the system selects “helpful” notes, which are displayed alongside the post for users to see.:
4. Generalized Model of Crowdsourced Context Systems
We coin the term Crowdsourced Context System to describe these systems, a novel and important addition to the content moderation toolbox. Such systems have several core characteristics. First, CCS rely on contributions crowdsourced from users of the platform. Second, these systems are designed to specifically elicit context from the crowd in order to help users understand and reflect on the veracity of the original post (as opposed to other types of responses, like, say, the most humorous reply). Third, these systems supplement user content by displaying additional text, or counterspeech (Peng and Grimmelmann, 2024), alongside the original posts. Finally, these systems are natively integrated into social media platforms. Unlike organic crowd behavior that emerges among users, these systems are built and operated by the platforms, giving platform designers the power to influence how the systems function.
Figure 2 shows a generalized model of a CCS. Such systems are composed of several core entities, which we denote with italics through the rest of the paper. Users create posts on the platform. Contributors then provide context, in the form of text or media snippets called notes, which are associated with specific posts. At first these notes are only visible to other contributors, who assess them and produce ratings of their helpfulness. The platform uses these ratings to identify “helpful” notes, which are then published and displayed alongside the associated post everywhere that it appears on the platform, as seen in Figure 1. In implementing this general model, platforms make many individual design decisions, which we will address later.
5. Methods
To better understand CCS and their design possibilities, we performed two distinct reviews. First, we performed a systematic literature review of existing research about CCS. Second, we performed inductive thematic analysis and affinity diagramming on a set of CCS design features extracted from documents describing the real-world CCS implementations of several major social media platforms.
5.1. Systematic Literature Review
Records identified through database searching (n = 205) Duplicates excluded (n = 8) Records after removing duplicates (n = 197) Records excluded due to title (n = 39) Records after title screen (n = 158) Records excluded due to abstract (n = 124) Records after abstract screen (n = 34) Reasons for full-text exclusion (n = 4) RQs are not about a CCS (n = 4) Studies included in qualitative synthesis (n = 30) IdentificationScreeningEligibilityIncluded
Figure 3. PRISMA flow diagram (Page et al., 2021) for our systematic literature review.:
In order to synthesize the research community’s understanding of Crowdsourced Context Systems, we performed a PRISMA-style systematic literature review (Page et al., 2021). Figure 3 shows a flow diagram summarizing our literature search and filtering process. Our goal was to identify original research that directly studies CCS. Since Twitter’s implementation was the first and most well-known CCS, we specifically used “Birdwatch” and “Community Notes” as keywords, as well as “Twitter” and “X”, with the expectation that any emerging systems would reference them. We additionally restricted our search to articles that have been written since 2021, when Birdwatch first launched. Note that since this is an emerging technology and the set of articles is still limited, we decided to include sources of knowledge that are valued in the HCI community but are sometimes excluded from literature reviews, such as pre-prints and non-full-length conference submissions. We thus settled on the following inclusion and exclusion criteria:
Inclusion criteria: the study is about Birdwatch or Community Notes or a similar system, either “in the wild” or in a controlled setting.
Exclusion criteria: The article was not written in English, does not present original research, is only an abstract/a summary of other research, or is not about the functioning of a CCS.
Search Strategy
We conducted our database search in July 2025. We searched seven databases that are well-known in the fields of HCI and social media research. We began with six that were used in another recent HCI literature review of social media research (Zhang et al., 2023): ACM Digital Library, IEEE Xplore, SAGE Journals, ScienceDirect (Elsevier), Taylor & Francis Online (Tandfonline), Emerald Insight. We additionally searched arXiv in order to also include preprints, which are common in these fields.
We searched the full-text, title, abstract, and metadata of articles using the following case-insensitive query string, and limit our results to articles published in 2021 or later:
("community notes" OR "birdwatch") AND ("x" OR "twitter")
This search resulted in n=205 results.
Title Screening
We screened the search results by first considering their title. One researcher read the title for each result and removed articles that were either duplicates or were obviously off-topic, such as those that were not about social media systems. This resulted in n=158 articles.
Abstract Screening
One researcher read the abstracts of the remaining articles to further screen the results. We removed redundant articles that only summarized research that was contained elsewhere in our sample, such as books and book chapters, editorials, abstracts, keynotes, and complete conference proceedings. We also removed articles that were topically irrelevant, such as those that did not focus on CCS, or that used data from a CCS, but only to answer research questions unrelated to the system. This resulted in n=34 articles.
Full-Text Screening
Two researchers then read the full-text of the remaining 34 articles and applied the inclusion and exclusion criteria in order to produce the final set of papers. The researchers met during the process to compare their results and discuss any differences or challenging cases until they were in agreement. This produced the final set of n=30 articles.
Data Extraction
Two researchers then agreed upon data to extract from each article in the final set, which included the data source, the research questions, the main findings, and any design implications mentioned. The two researchers independently read the final set of articles and extracted this information. They then compared values, discussed any areas of disagreement, and came to a consensus for the extracted values for each paper.
Next the two researchers independently grouped like-papers into affinity groups (Beyer and Holtzblatt, 1999) based on the focus of the study. The researchers reconvened, compared their groups, and discussed the reasons for their groupings. Several papers were shifted between groupings based on the discussion. Once the researchers agreed upon the groupings, they came up with labels to describe each group.
5.2. Design Feature Identification
In order to identify the design decisions that shape CCS, two authors analyzed four real-world implementations of CCS: Twitter/X’s Community Notes [^5] https://x.com/i/communitynotes/, Meta’s Community Notes [^6] https://transparency.meta.com/features/community-notes (which covers Facebook, Instagram, and Threads), TikTok’s Footnotes [^7] https://support.TikTok.com/en/using-TikTok/exploring-videos/footnotes, and YouTube’s Notes [^8] https://support.google.com/YouTube/answer/14925346?hl=en. Our analysis is based on public, company-written documentation and communications about the systems and their features, which we will collectively refer to as articles. In addition, we had access to some of the systems, which we used to complement our understanding. Unfortunately, while we attempted to become Contributors for all of the platforms, some (TikTok, YouTube) did not grant us access.
The researchers independently read through the articles one at a time and extracted design details into a spreadsheet, which resulted in a set of “design features” for each platform. After independently reviewing the articles for a platform, the researchers convened and performed a round of affinity diagramming (Beyer and Holtzblatt, 1999), in which they wrote the design features on sticky notes and placed them on a whiteboard. They placed “similar” design features together, and through this inductive method identified “feature groups”. The researchers then performed thematic analysis (Braun and Clarke, 2006) to identify the themes of the groupings and came up with provisional group names. They repeated this process with the articles for each platform, reconvening to do another round of affinity diagramming after each round of feature extraction. These design features and feature groups were further synthesized into the aspects, dimensions, and options of the design space, which we present below in Section 7.
6. What Does the Research Community Know About CCS?
Study Focus Primary Method Sample Paper
System Impact (n=10) Data Analysis: X Community Notes Data Diffusion of Community Fact-Checked Misinformation on Twitter (Drolsbach and Pröllochs (2023))
Contributor Behavior (n=9) Data Analysis: X Community Notes Data Birds of a feather don’t fact-check each other: Partisanship and the evaluation of news in Twitter’s Birdwatch crowdsourced fact-checking program (Allen et al. (2022))
User’s Perceptions (n=6) User studies Checking the Fact-Checkers: The Role of Source Type, Perceived Credibility, and Individual Differences in Fact-Checking Effectiveness (Liu et al. (2023))
System Design (n=5) User studies, design specs, algorithm analysis Supernotes: Driving Consensus in Crowd-Sourced Fact-Checking (De et al. (2025))
Table 1. Summary of the existing research found in our literature review.:
Our systematic literature review identified n=30 papers about Crowdsourced Context Systems, summarized in Table 1. These papers fall into several high-level categories: studies of the system’s impact, contributor behavior, users’ perceptions of crowdsourced context, and alternative system designs. Some studies contained elements of multiple categories, in which case we grouped them based on their most significant contribution.
Many of the published CCS research studies focused on the CCS’s impact on the platform where it is implemented (n=10 papers). These studies all performed analysis of data from X’s Community Notes. These studies asked both causal questions, such as about the effect that displaying a note has on post diffusion (Drolsbach and Pröllochs, 2023; Chuai et al., 2024b), as well as descriptive questions about the performance of X’s system in several specific use cases (Chujyo et al., 2025; Bouchaud and Ramaciotti, 2025). The studies found strong evidence that CCS can both reduce the spread (Drolsbach and Pröllochs, 2023) and believability (Wojcik et al., 2022) of misleading posts. The research also argued that the bridging-based algorithm in X’s Community Notes system has demonstrated it can identify high-quality crowdsourced notes that are likely to be perceived as helpful by users across an ideological spectrum (Bouchaud and Ramaciotti, 2025). However, the studies also showed that this bridging approach takes a significant amount of time to identify “helpful” notes for misleading posts. As a result, there is often a significant delay before notes are displayed to users (Renault et al., 2024; Chuai et al., 2024a, b). Additionally, studies have shown that the vast majority of suggested notes ($\sim 90\%$) are never published due to either a lack of ratings or consensus  (Chujyo et al., 2025; Wirtschafter and Majumder, 2023).
Another common type of research study (n=9) examined CCS contributor behavior. While one of these studies used interviews as a data source (Yoon et al., 2025), all of the others relied on the open data from X’s Community Notes. The studies asked descriptive questions about how contributors to the system behaved, such as what types of posts (Pilarski et al., 2023; Saeed et al., 2022) and accounts (Allen et al., 2022) contributors tend to leave notes on, and what type of content contributors tend to include in notes (Jones et al., 2022; Kuuse et al., 2025). Two key findings from these studies are that 1) the crowd exerts political bias in the posts that they select to write notes for and the ratings that they assign to notes (Allen et al., 2022); and 2) the crowd behaves differently than professional fact-checkers (Pilarski et al., 2023), yet often links to them as sources (Borenstein et al., 2025).
A third group of studies sought to understand user’s perceptions of CCS (n=6). These studies used experiments (Liu et al., 2023), interviews (Shusas and Forte, 2024), and data analysis (Hameleers, 2024) of X’s data to answer descriptive questions about how individuals perceived the credibility and helpfulness of context provided by CCS. These studies have demonstrated that certain types of notes are perceived as more helpful than others, such as those that support their claims with links to unbiased (i.e., assessed as politically neutral by a third-party organization [^9] https://mediabiasfactcheck.com/) external sources (Solovev and Pröllochs, 2025). There is also evidence that users perceive notes from CCS as just as, or even more, credible than those from professional fact checkers (Liu et al., 2023; Shusas and Forte, 2024).
The final group of studies were design studies that either proposed or built a variation of a CCS (n=5). These papers focused on ways to improve the quantity (Li et al., 2025) and quality (Mohammadi and Yasseri, 2025) of notes, as well as techniques to improve the performance of the note curation algorithm (Mujumdar and Kumar, 2022). The bridging-based algorithm for selecting “helpful” notes was introduced in one of these papers (Wojcik et al., 2022), as a solution to problems of bias (Allen et al., 2022) in a majority-rule note selection system. Recent papers have proposed ways of incorporating AI into the note-writing process and found that LLMs have the potential to generate helpful notes (De et al., 2025; Mohammadi and Yasseri, 2025). While these studies begin to imagine alternative system designs, much is still unexplored. As we show next, the potential design space is much wider than what is covered in these studies.
7. A Design Space for CCS
Our analysis of the various CCS implementations revealed areas of common characteristics shared between the systems as well as points for variation. Analyzing these features allows us to construct a design space of possibilities. Within this design space, we identify six key aspects of CCS:
  1. (1)
    Participation: Who uses the system and what are different users are allowed to do?
  2. (2)
    Inputs: What shape do contributions to the system take?
  3. (3)
    Curation: How are suggested notes selected to be displayed?
  4. (4)
    Presentation: How are selected notes displayed?
  5. (5)
    Platform Treatment: How does the platform treat posts that have displayed notes?
  6. (6)
    Transparency: How can the public observe and evaluate the functioning of the system?
Within each aspect, we identify several dimensions (i.e., points of potential variation), which we frame as design questions, and options (i.e., potential answers for each dimension) based on the features that we observed in our analysis. In the tables in the sections that follow, we indicate when we found evidence that a platform implemented a certain option. However, platforms did not always document all of their design choices, so our tables may be incomplete. In constructing the design space, we additionally include several speculative options that we did not see in any current implementations. The sections that follow detail the dimensions and options for each aspect.
7.1. Participation
Dimension Implementing Platforms
Who sees published notes?
All users. Meta, Twitter/X
Only users from a specific geo/market. TikTok, YouTube
Only users that have opted in.
Who are the context contributors?
Users. TikTok, Meta, Twitter/X, YouTube
Platform employees/contractors.
Third parties.
AI systems. Twitter/X
Who are the rating contributors?
Users. TikTok, Meta, Twitter/X
Platform employees/contractors. YouTube
Third parties.
AI Systems.
What must a user do to become a context contributor?
Nothing: all users are contributors by default.
Nothing: users cannot be context contributors. YouTube
Apply. TikTok, Meta, Twitter/X
Receive an invite from the platform. YouTube
Be admitted by the platform. TikTok, Meta, Twitter/X
Meet a quality threshold of past performance. Twitter/X
What must a user do to become a rating contributor?
Nothing: all users are contributors by default.
Nothing: users cannot be rating contributors.
Apply. TikTok, Meta, Twitter/X, YouTube
Receive an invite from the platform. YouTube
Be admitted by the platform. TikTok, Meta, Twitter/X, YouTube
Meet a quality threshold of past performance.
What types of eligibility requirements limit who can be a contributor?
User-age requirements. TikTok, Meta
Account-age requirements. TikTok, Meta, Twitter/X, YouTube
Lack of policy-violations. TikTok, Meta, Twitter/X, YouTube
Geographic requirements. TikTok, Meta, Twitter/X
Language requirements. Meta, YouTube
Account authentication requirements. TikTok, Meta, Twitter/X
Proof of some real-world credentials.
Is the number of contributors limited?
Yes: platform has discretion over who they admit. TikTok, Meta, Twitter/X, YouTube
No: all who meet eligibility criteria are automatically admitted.
How are contributors incentivized?
Paid by the platform. YouTube
Pseudonymous reputation. Twitter/X
Account reputation.
Unclear: contributor behavior is not visible.
How are notes and ratings solicited?
Contributors can add notes to posts they organically come across while using the platform. TikTok, Meta, Twitter/X, YouTube
Contributors have feeds of posts that need notes, and notes that need ratings. Meta, Twitter/X
Platform sends notifications to contributors requesting review of specific posts and notes. Twitter/X
Users can request notes on posts or additional review on published notes. Meta, Twitter/X, YouTube
Table 2. Participation dimensions and options, as well as platforms that have made this design choice.:
We define participation as the design choices that shape who uses the system and what different users are allowed to do. Table 2 shows all of the participation dimensions and options. Choices that platforms make in this aspect range from those that encourage user participation in the system, to those that discourage it.
The most basic dimension is, “Who sees published Notes?” The options that platforms choose as answers to this question will greatly affect the system’s impact. The most participatory implementations like X’s allows all platform users to see these notes, while less participatory implementations limit the audience in some way beyond a user’s control, for example by their location. An even less participatory option would be to make notes an optional feature that users must opt in to see, which would greatly limit the system’s impact.
Other dimensions of participation determine who can contribute to the system, such as: “Who are the context contributors?” These dimensions shape the quality, quantity, and variety of the notes that the system produces. All implementations we examined rely on platform users as contributors, but we see some variation in what these users are allowed to contribute. For example, YouTube currently does not allow users to provide ratings; instead, those are provided by third-parties who are paid by the platform. X allows AI systems to contribute notes via an API. All of the implementations impose eligibility restrictions on users, ensuring that platforms maintain discretion over who is admitted as a contributor. Yet we can imagine a system that would automatically admit all eligible applicants, or even one with no eligibility restrictions, where all users would be contributors by default.
Besides decisions that limit the set of possible contributors, it is also important to ask questions about how platforms encourage these contributors to participate, such as: “How are contributors incentivized?” While it would be possible for platforms to pay contributors, all of the implementations that we examined rely on volunteer labor to some extent. These approaches may choose to incentivize contributions through reputation systems, for example by publicly associating contributors with either their user accounts or with pseudonyms. There is also the important question of: “How are notes and ratings solicited?” All platforms allowed contributors to add notes to posts that they organically come across while using the platform. However, this design choice could result in skewed post selection if, for example, contributors follow accounts from one end of the ideological spectrum more than the other. Some platforms may also take more active approaches that solicit notes for specific kinds of posts, or ratings for specific suggested notes, either through in-app feeds or notifications, as shown in Table 2. These approaches could be used to counter bias in the system, or to encourage notes on posts that are most likely to do harm, such as those that are going viral.
7.2. Inputs
Dimension Implementing Platforms
What tools are note contributors given?
Free-text inputs TikTok, Meta, Twitter/X, YouTube
Ability to add source links. Meta, YouTube
Ability to add references to a portion of the post (e.g. timestamp in the video being noted). YouTube
AI note writing assistance.
API for AI note creation. Twitter/X
What tools are rating contributors given?
A fixed set of options (i.e., helpful, somewhat helpful, not helpful) to select from. TikTok, Meta, Twitter/X, YouTube
A free-text explanation field. YouTube
A fixed set of explanation options. Meta, Twitter/X
API for AI rating creation.
Which posts can have notes left on them?
All. Twitter/X
Posts by specific account types (public, verified, pages). TikTok, Meta
Posts of specific types of content (no ads, none that may feature minors). YouTube
Posts that reach a certain level of engagement.
Posts that have been flagged by users as needing additional context.
What restrictions apply to notes?
Must adhere to community guidelines. TikTok, Meta, Twitter/X, YouTube
Length restrictions. Meta, YouTube
Cannot be edited once submitted. Meta, Twitter/X
Cannot be deleted once submitted.
Source URL required. Meta
What restrictions apply to ratings?
Must adhere to community guidelines. YouTube
Length restrictions. YouTube
Cannot be edited once submitted. Twitter/X, YouTube
Cannot be deleted once submitted. Twitter/X, YouTube
Can the published note on a post change over time?
Published notes can be hidden by the crowd indefinitely. TikTok, Meta, YouTube
Note authors can edit or delete their notes. Meta, YouTube
Post authors can edit/remove notes on their posts.
Published notes are ”locked” after a certain amount of time. Twitter/X
Table 3. Inputs dimensions and options, as well as platforms that have made this design choice.:
We define inputs as the design choices that shape the type and content of contributions to the system. Table 3 shows the associated dimensions and options for the inputs dimension.
Contributions are shaped by the affordances that platforms provide to contributors to create their input, including the user interfaces for producing notes and ratings. These interfaces may allow free-form, or structured, input. Such decisions may have an effect on what information contributions contain and what they look like. For example, Meta’s implementations have a required “source URL” field, which based on the findings mentioned in Section 6, may lead to more helpful notes. The content of contributions is also shaped by additional restrictions placed on them by platforms. For example, all of the implementations that we examined required contributions to adhere to the platform’s content policy. Finally, design choices that require contributors select from a set of fixed options, may be both easier to use and produce more consistent inputs. In these ways, design decisions in this aspect will affect the volume and quality of both notes and ratings.
A final interesting question is, “Can the published note on a post change over time?”. Some implementations will give contributors the ability to edit or delete their notes, while others will prevent this. Additionally, the published note on a post can change if the curation process that selects notes (discussed below) is re-run over time. To deal with this, X’s implementation added a feature that “locks” the published note after a certain amount of time, preventing it from changing.
7.3. Curation
Dimension Implementing Platforms
What determines which notes are shown alongside posts?
Algorithmic decision based on data from contributors. TikTok, Meta, Twitter/X, YouTube
Editorial decision made by the platform.
Editorial decision made by contributors.
What (alleged) control does the platform have over what is displayed?
Platform chooses what is or is not published.
Platform can edit note content.
Platform can remove notes if they violate platform content policy. TikTok, Meta, Twitter/X, YouTube
Platform has no control.
What data is used to make an algorithmic decision?
Ratings. Twitter/X
Note content (length, inclusion of a link, etc).
A note author’s past platform behavior as a contributor.
A note author’s past platform behavior as a regular User.
Demographic/personal information about the note author.
Rating contributor’s past platform behavior as a contributor. Twitter/X
Rating contributor’s past platform behavior as a regular user.
Demographic/personal information about the rating contributor.
What is the aim of the algorithmic decision?
Appeal across the ideological spectrum (Bridging). TikTok, Meta, Twitter/X, YouTube
Other?
Table 4. Curation dimensions and options, as well as platforms that have made this design choice.:
We define curation as the design choices that affect how notes are selected to be displayed. Table 4 shows all of the associated dimensions and options. Design decisions here affect the quality of the notes that are displayed, the potential for bias in the note selection, as well as the speed and volume at which they appear.
Perhaps the most important design question is “What process determines which notes are shown alongside posts?” The existing systems that have recently grown in popularity all involve algorithmic decisions by bridging-based algorithms that, broadly speaking, identify notes that have been rated as “helpful” by contributors who have disagreed in their past ratings. Even though these bridging algorithms are currently popular, alternatives to this scheme are possible: for example, Twitter’s first implementation of Birdwatch used a majority voting algorithm. Alternatively, it could be possible to achieve bridging goals using contributor information other than their past ratings. Algorithmic approaches are also not the only option, as one could imagine a CCS that selects notes via editorial decisions by either the platform or contributors. Another possible approach is to select notes via digital juries (Fan and Zhang, 2020) of users, which could be selected either randomly or based on other criteria.
A related question is, “What (alleged) control does the platform have over what is displayed?” Platforms can exert varying levels of control over the output of these systems, although current implementations all intentionally minimize the platform’s role. Still, all platforms do retain control to make editorial decisions in the case of content policy violations.
7.4. Presentation
Dimension Implementing Platforms
Where are published notes presented to users?
Below the text of the original post. TikTok, Meta, Twitter/X, YouTube
Above the text of the original post.
In an overlay on top of the original post.
On both the original post and any re-shares of it. Twitter/X
Who is notified about notes when they are published?
Original post authors. TikTok, Meta, Twitter/X, YouTube
Contributors who have written/rated the published note. Meta, YouTube
Every user that replied, liked, or reposted the original post. Meta, Twitter/X
Every user that was exposed to the original post.
How many notes are shown on a post?
At most one. TikTok, Meta, Twitter/X, YouTube
Multiple.
How is the author of a note shown on a post?
The author of the note is anonymous/omitted completely. TikTok, Meta, Twitter/X, YouTube
The author of the note is pseudonymous. Twitter/X
The actual identity of the note author is shown.
How are a note’s ratings shown on a post?
Only whether or not it is “helpful”. TikTok, Meta, Twitter/X, YouTube
Some indicator of how helpful it is.
Table 5. Presentation dimensions and options, as well as platforms that have made this design choice.:
We define presentation as the design choices about how published notes are displayed to users. Table 5 shows all of the associated dimensions and options. Design decisions here affect the impact of a published note on users who see the associated post.
The biggest question is, “Where are published notes presented to users?” Answers to this question affect whether or not users will ever see published notes. Note presentation can range from aggressive to subtle: notes can be mandatory overlays that users must dismiss to see the associated post; notes can be shown alongside posts where they can be easily ignored; or platforms may simply indicate that a post has a note and require users to click to view them. Additionally, there is the question of which instances of the post have the published note displayed: just the original, or also re-shares of the post that appear across the platform. This will affect the number of users who have the potential of being informed by a published note. Platforms must also decide: “Who is notified about notes when they are published?” Notifications here can be used to retroactively influence the understanding of users who saw a post before it had a published note.
Other important questions affect the amount of information about notes that is conveyed to users. Twitter’s original implementation of Birdwatch showed several “helpful” notes at a time, but all of the current implementations that we examined only show one. Relatedly, platforms may provide details about a note’s author or ratings, which could inform how a user interprets a note.
7.5. Platform Treatment
Dimension Implementing Platforms
How are posts with notes on them treated by the platform?
No different from any other posts. TikTok, Meta, Twitter/X, YouTube
Limit monetization options.
Limit positioning next to ads. Meta
Limit spread (disabling reshares, deranking in feeds)
Increase spread (boost ranking in feeds)
How do notes interact with other content moderation features?
Replaces third-party fact checking program. Meta, Twitter/X
Complements platform-provided content moderation. TikTok, Meta, Twitter/X, YouTube
Notes are presented on posts alongside other “information reliability” indicators. TikTok, YouTube
Table 6. Platform Treatment dimensions and options, as well as platforms that have made this design choice.:
We define Platform Treatment as the design choices about how the platform treats posts that have published notes. Table 6 shows all of the associated dimensions and options. Design decisions here affect the impact of CCS systems on the spread of posts.
The central question is, “How are posts with notes on them treated by the platform?” All current implementations make a point of emphasizing that they do not algorithmically penalize posts with notes on them (for example, by down-ranking in recommendation algorithms). But this need not be the case: perhaps certain types of notes provide a useful signal about post quality and platforms may want to suppress them. Alternatively, the platform may impose restrictions on posts with notes, like prohibiting them from being monetized or from displaying next to ads.
It is also important to ask “How do notes interact with other content moderation features?” Some platforms have shut down their fact-checking programs and replaced them with CCS. But, all of the platforms that we surveyed still have content moderation systems of some kind, which a CCS complements. Finally, YouTube and TikTok display multiple “information reliability” indicators alongside posts in tandem with notes, such as platform-provided banners or labels.
7.6. Transparency
Dimension Implementing Platforms
How public is the note curation system?
Open source. Twitter/X
Open data. Twitter/X
Fully verifiable.
Not public. TikTok, Meta, YouTube
How public is the contributor admission process?
Open source.
Open data.
Fully verifiable.
Not public. TikTok, Meta, Twitter/X, YouTube
How public is the data needed to evaluate the impact of notes on engagement?
Open data.
Available to researchers.
Available to paying customers. Twitter/X
Not public. TikTok, Meta, YouTube
What sort of documentation about the system is public?
Blog posts. TikTok, Meta, YouTube
Help articles/FAQs. Meta
Technical documentation. Twitter/X
None.
Table 7. Transparency dimensions and options, as well as platforms that have made this design choice.:
We define transparency as the design choices that influence how the public can observe and evaluate the functioning of the system. Table 7 shows all of the associated dimensions and options. Design decisions here may affect the perception of the CCS by users, the public, and the research community.
Of central interest is, “How public is the note curation system?” The system’s source code and data can be public in order to allow outside study. In addition to being public, platforms could take additional steps to make their source code and data verifiable, or to allow outsiders to verify that the public code and data is indeed the same code and data that is powering the platform’s implementation. Or, platforms could do none of this and instead keep everything private, which is the approach chosen by all platforms except for X.
To fully evaluate the efficacy of such a system, we also need to ask, “How public is the data needed to evaluate the impact of notes on engagement?” This data would allow the public to holistically evaluate the CCS as opposed to other systems of content moderation. Currently this data is very hard to access, which makes holistically assessing the efficacy of these systems a challenge.
8. Normative Concerns
CCS are complex sociotechnical systems (Cherns, 1976) that integrate people, algorithms, and speech. These systems’ growing influence over the information shown on major social media platforms means that the design choices made in their implementation can have significant social consequences (Friedman and Nissenbaum, 1996; Flanagan et al., 2008). With such high stakes, it is crucial that a human-centered research agenda consider the implications of different design choices, in order to holistically evaluate the impacts of these novel systems.
The design space that we present in Section  7 highlights some of the crucial choices that are made when designing CCS. How are we to weigh the merits of different designs? One approach is to identify the primary normative concerns of such a system. We emphasize three interconnected normative concerns that should guide CCS design: user informedness, distribution of power, and fairness.
8.1. User Informedness
While social media serves many different purposes, CCS are specifically concerned with social media’s ability to inform the public. Such systems should thus be evaluated by how well they facilitate user informedness: for example, the degree to which they inform users about the accuracy of information, minimize the spread of misleading posts, or otherwise improve information quality on a platform. Our literature review surfaced ample evidence that the notes produced by CCS do have the power to increase user informedness, both by decreasing an individual’s belief in misleading information, and by decreasing the spread of this information (Wojcik et al., 2022; Renault et al., 2024; Slaughter et al., 2025; Chuai et al., 2024a, b; Drolsbach and Pröllochs, 2023). But while published notes have been shown to be effective at promoting understanding, we must also question how effective CCS are at producing these notes. Several of these same studies found that Twitter’s implementation of CCS produces notes too slowly to prevent the spread of the most viral content (Chuai et al., 2024b, a; Renault et al., 2024). There is also a significant gap in the research on the question of coverage, or how many of the misleading posts on a platform receive “helpful” notes, which is currently unanswered. The research suggests that there is a CCS design trade-off between speed/coverage and quality/helpfulness: X’s implementation has optimized for the latter, perhaps at the expense of the former. Whether this is the right decision should be up for debate.
For designers of CCS hoping to improve the system’s speed and coverage, there are several aspects from our design space that could be relevant. For example, within the curation aspect, designers could explore bridging algorithm changes that could lead to faster identification of “helpful” notes. Additionally, design changes within the participation and inputs aspects, such as those that would allow a larger contributor base or that would actively solicit notes on the most viral posts, could lead to a higher volume of notes and ratings, and thus quicker selection of helpful notes. When it comes to the difficulty of assessing the coverage of CCS, design changes within the transparency aspect could make this easier to study, perhaps by providing easier access to post engagement metrics.
8.2. Distribution of Power
One of the selling points of CCS is that they shift power over content moderation from the platform to the community. Giving users more of a say over the content that they see on their platform can make moderation decisions more legitimate and trusted (Shusas and Forte, 2024; Liu et al., 2023; Drolsbach et al., 2024), which is an admirable goal. Yet such claims require further scrutiny, and CCS should be evaluated to the extent that they actually empower users and do not merely allow platforms to shirk responsibility for misleading content. Evidence from our literature review suggests that bridging-based algorithms, by design, fail to produce notes for the most polarizing content (Bouchaud and Ramaciotti, 2025). For example, X’s implementation of the bridging algorithm requires polarized contributors to agree that a note is helpful in order to publish it. The editorial function of the bridging algorithm can thus limit the type and scope of contributions in two important ways: in terms of framing, the algorithm may prefer centrist points of view and content that could be agreeable to both sides of an issue; in terms of selection, agreement on a note may not occur for contentious topics, preventing its publication. Whether or not users at the ends of the ideological spectrum should be empowered to make moderation decisions is a valid question, but it is uncontroversial that the most polarizing content needs moderation, too. Research has found evidence of harmful content being handled poorly by the system, either because potential notes never reach helpful status (Chujyo et al., 2025), or because notes that do are not actually helpful for certain types of content, such as disinformation that is couched in humor (Matamoros-Fernández and Jude, 2025). Zooming out, the low publishing rate for suggested notes (around  10% on X (Wirtschafter and Majumder, 2023) and  6% on Meta [^10] https://x.com/guyro/status/1965800300557857053) gives us pause that the system is actually empowering these note authors. A related, underexplored area for research involves studying which note contributors are empowered by CCS. This research area is difficult to study holistically, as these systems are pseudonymous, but some work has begun to explore it qualitatively (Yoon et al., 2025) and via partnerships with platforms (Allen et al., 2022). More research is needed to understand if CCS empower users of certain demographics or political views over others.
Our design space offers several ways to further empower users via CCS. First and foremost is innovation in the participation aspect that would increase the set and diversity of contributors. The challenge for the platforms will be ensuing that quality remains as the participating contributor base increases, which is a tension that has been thoroughly studied in the crowdsourcing literature (Daniel et al., 2018), for example. Innovation in the inputs aspect may increase the quality of notes across the board, perhaps by giving note authors helpful feedback (Mohammadi and Yasseri, 2025), empowering more users to make successful contributions to the system. Another route to shift power to users is to disempower the platform. Though CCS are often described as ways to reduce platforms’ centrality in content moderation, platforms still exert power over the system in important ways, such as by deciding who is admitted as a contributor. Platforms could further guarantee their neutrality through changes to the transparency aspect, such as by making their contributor selection process public. But considering that X’s CCS implementation is the only even partially open system that we examined, a first step is for other platforms to follow X’s lead and open source their note curation algorithm and the crowdsourced data that they collect.
8.3. Fairness
CCS are often touted for their potential to implement content moderation in an unbiased way. Yet these systems are not inherently free from bias. Results from our literature review detail bias in contributor behavior, in terms of the types of accounts that they leave notes on, as well as how they rate suggested notes (Allen et al., 2022). Bridging algorithms have helped mitigate bias in the note curation process, with the caveats noted above, but additional bias can still be introduced—for example, in the note creation process, or in the selection of contributors. To give a concrete example: X’s bridging algorithm finds agreement among contributors, not the entire user base. If contributors’ beliefs are not representative of users’ beliefs then the algorithm’s outcomes may be a biased representation of the community, and not necessarily have broad appeal among users. There is some evidence that so far X’s implementation has not suffered from this problem, but it is worth considering. Also, accusations of bias in top-down moderation are often due to the opacity of such approaches and may be mitigated by transparency efforts. Only X’s implementation has provided any transparency, and yet the lack of transparency about the contributor base has made analyzing them for bias impossible. Finally, the approach of displaying context rather than removing posts is arguably fairer in that it lessens the damage from false positives, by empowering users to come to their own conclusions about content’s veracity (Peng and Grimmelmann, 2024). However, there is a trade off: leaving up questionable content, even with helpful context, can potentially do harm to those targeted by the content.
Several aspects from our design space can influence the fairness properties of a CCS. First off, all platforms have taken the approach that it is fairer not to remove misleading content. However designers may explore other options from the platform treatment dimension to identify certain cases where the platform should take action on content flagged by the crowd as misleading. Second, current implementations tend to display a single, definitive note, even if there may be several that are determined to be “helpful”. The presentation aspect offers alternatives that could further the aim of allowing users to come to their own conclusions, such as by showing multiple notes when there are multiple with helpful ratings. Additionally showing information about the note author could be helpful, as it could help users identify the credibility of the source. This may be at odds with contributor privacy, which current implementations have opted to protect. Finally, innovations in the curation aspect could explore other methods of selecting notes to display that may be regarded as fairer. Current implementations of the bridging algorithm use past ratings to identify contributors with diverse perspectives: perhaps it would be fairer to instead use in-app behavior? This could provide a fuller picture of a contributor’s ideological leanings. Encouraging more transparency would also give assurances of fairness, both in how contributors are selected, and how notes are published.
9. Discussion
Through our systematic literature review and analysis of real-world CSS implementations, we have produced a theoretical framework that can help the HCI research community better reason about, and design, CCS. These systems were created with the stated goal of addressing the misinformation problem (Coleman, 2021), but it is unclear if the platforms implementing them are serious about this aim. For example, many see Meta’s choice to replace their U.S. fact-checking program with a CCS (Kaplan, 2025) as a thinly-veiled effort to curry political favor with a new presidential administration that is critical of fact-checking (Nix et al., 2025). In fact, from a platform’s perspective, CCS have other appealing properties: they cost less than fleets of fact-checkers, they shield platforms from accusations of biased policy enforcement, and they offload the difficult responsibility of distinguishing fact from fiction onto users. It may be expedient for platforms to build CCS in the current moment, but it is unclear if such systems will survive if the political climate shifts. Still, as long as CCS are in a position to exert such influence over the information ecosystem, it is important for the HCI research community to study them. Most importantly, we emphasize the need to better understand how CCS perform compared to other content moderation systems. CCS may perform well in certain cases, but if on the whole they perform worse than the systems they are replacing, for example on the question of coverage, they should be used carefully.
While it has been suggested that they can, CCS cannot effectively replace a fuller spectrum of content moderation tools, and were not designed to do so (Elliott, [n. d.]). Content moderation is a broad term that encompasses a variety of different goals, including preventing crimes, mitigating lawful but awful behavior, and promoting high quality content. The tools used as a part of content moderation are similarly varied and include content take-downs, labeling, and account bans. Systems that provide context are much narrower in the scope of their goal and mechanism: they aim to better inform users by providing them with additional information. And yet, the empirical results summarized in this paper demonstrate that CCS may play a meaningful role in the content moderation ecosystem, and it is worth exploring how they may integrate with other parts of it. For example, the research that we reviewed found that notes often build on the work done by professional fact-checkers (Borenstein et al., 2025). Might this relationship flow the other way, as well? Perhaps data from CCS systems could be used by professional fact-checkers to cover a wider range of issues. Or, the ability of CCS systems to identify notes with broad appeal could help professionals craft their own fact-checks that do the same.
Our design space reveals several ways forward for a HCI research agenda on CCS. Several aspects are underexplored by current research, particularly the Participation aspect. Encouraging positive participation in collaborative computing systems is a classic topic in HCI research (Kraut and Resnick, 2012) and one in which the community could make strong contributions. One dimension of participation which has not yet been explored is how platforms may play a more active role in soliciting notes and ratings from contributors. Past HCI work on how to best make use of the various skills of crowds may be instructive here (Bernstein et al., 2015; Kittur et al., 2013). It is worth noting that the more consequential these systems become, the more they will be targets for negative forms of participation as well, primarily gaming (Daniel et al., 2018; Epstein et al., 2020). Bridging algorithms are designed to be resistant to certain types of gaming (Wojcik et al., 2022), but vulnerabilities may emerge. One other area for study is the Presentation aspect, specifically how different approaches to visually displaying notes may affect user understanding of misleading claims (Gamage et al., 2025; Martel and Rand, 2023). Another promising area is the Inputs aspect, where recent work has begun to explore the ways in which LLMs can supplement human effort and increase the quality and quantity of notes (De et al., 2025; Mohammadi and Yasseri, 2025). Of particular interest is X’s recent decision to allow AI note contributors via a new API (Li et al., 2025). How such a decision affects the quantity and quality of notes on the platform should be a near-term focus of HCI research. Of course, studying how changes to CCS impact information dynamics across a platform will be difficult to study without more innovations in the Transparency aspect. X’s approach to transparency in their Community Notes system has been a boon for for studying CCS. Hopefully other platforms will soon adopt similar approaches to support ongoing studies of CCS effectiveness.
10. Conclusion
Crowdsourced Context Systems are an increasingly prevalent part of the content moderation strategies of many prominent platforms. To understand these systems, their impacts, and how they may be better designed, we conducted a systematic literature review of existing research studying implementations of CCS and analyzed several major social media platforms’ implementations of such systems. Despite recent trends towards the adoption of these systems, we emphasize that they are only one piece of a larger content moderation suite, and suggest against using them as drop-in replacements for traditional fact-checking approaches. The framework we produced can be used to reason about these systems, and consists of a theoretical model, a design space, and several key normative concerns. Such systems have potential to address the challenge of misinformation, but to realize this potential, future design research from the HCI community must focus on ways to understand and improve these systems’ performance.