### v2

# Beyond Community Notes: A Framework for Understanding and Building Crowdsourced Context Systems for Social Media

Abstract.
Social media platforms are increasingly developing features that display crowdsourced context alongside posts, modeled after X’s Community Notes. These systems, which we term Crowdsourced Context Systems (CCS), have the potential to reshape our information ecosystem as major platforms embrace them as alternatives to top-down fact-checking. To deeply understand the features and implications of such systems, we perform a systematic literature review of existing CCS research and analyze several real-world CCS implementations. Based on our analysis, we develop a framework with two distinct components. First, we present a theoretical model to help conceptualize and define CCS. Second, we identify a design space encompassing six key aspects of CCS: participation, inputs, curation, presentation, platform treatment, and transparency. We discuss key normative implications of different CCS design and implementation choices. Our paper integrates these theoretical, design, and ethical perspectives to establish a foundation for future human-centered research on Crowdsourced Context Systems.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
Refer to caption Refer to caption :
Figure 1. Screenshots of a Community Note displayed on Instagram (left) and X (right).
1. Introduction
Social media platforms play an important role as sites of entertainment (herderDigitalJunkfoodSocial2024), news consumption (chouakiWhatNewsPeople2024a), and interpersonal connection (preeceEmpathicCommunitiesBalancing1999). At the same time, these platforms must contend with problems stemming from harmful user-uploaded content, which can take the form of abuse, spam, or misinformation (scheuermanFrameworkSeverityHarmful2021). Specifically, platforms have tried various forms of content moderation (grimmelmann2015virtues; lampeSlashdotBurnDistributed2004; jhaverHumanMachineCollaborationContent2019) to deal with misinformation, including centralized fact-checking efforts to label or remove it (gillespieCustodiansInternetPlatforms2018), as well as distributed approaches where users are empowered to tackle these challenges themselves (seeringReconsideringSelfModerationRole2020). While these approaches have had some success, they have also evoked backlash and accusations of bias from both users and lawmakers, particularly on the political right (jiangReasoningPoliticalBias2020; johnsonFactsAllegationsPolitical2023). Recently, a new approach has emerged: beginning with X/Twitter, major social media platforms are increasingly introducing systems that crowdsource “helpful context” to display alongside potentially problematic posts (IntroducingBirdwatchCommunitybaseda) (e.g. Figure 1). These systems, which we term Crowdsourced Context Systems (CCS), have seen growing adoption and impact, with some of the world’s largest platforms, including Meta (Facebook, Instagram, and Threads), TikTok, and YouTube all recently launching their own implementations. Meta has gone one step further, announcing that their CCS will replace their fact-checking program in the U.S. (heatheraMoreSpeechFewer2025).
As CCS adoption grows, these systems have the potential to reshape our information ecosystem. Early studies into such systems have produced promising results, finding that CCS can increase user informedness and decrease the spread of misinformation (wojcikBirdwatchCrowdWisdom2022b; rollout21; diffusion21), although it is still unclear whether CCS can effectively replace centralized fact-checking, as some platforms have done or may seek to do (heatheraMoreSpeechFewer2025). As CCS develop, novel implementation questions also come to the fore, including how to design these systems most effectively—for example, to provide reliable context quickly enough to prevent the spread of the most viral content (rollout21; chuaiCommunitybasedFactcheckingReduces2024a; renaultCollaborativelyAddingContext2024a). At this critical moment of change in the information ecosystem, it is important to understand how CCS function, how we should evaluate them, and how these systems may be better designed.
This paper makes two contributions towards the goal of better understanding CCS. Our first contribution is a systematic literature review of existing research studying CCS implementations. Our review finds significant research into CCS contributor behavior and the impact of displaying notes. However, the review highlights the fact that certain aspects of these systems are underexplored by existing research. We thus set out to identify additional facets of these systems that may affect their impact. Our second contribution is an analysis and comparison of real-world implementations of CCS by several major social media platforms, in order to create a framework to guide future HCI research on these novel and important systems. Our framework is composed of two parts. First, we offer a theoretical model of a CCS that identifies these systems’ core characteristics and what sets them apart from prior crowdsourced systems. Our model posits that while CCS use familiar moderation tools like crowdsourced information and context labels, they are unique in how they are natively integrated into platforms with the explicit goal of facilitating content interpretation. Second, we outline a design space (macleanQuestionsOptionsCriteria1996) for CCS in order to taxonomize the important design decisions made in their implementations. Design spaces are a conceptual tool used in HCI research that can guide work on complex socioetechnical systems by structuring the space of design possibilities (zhangFormFromDesignSpace2024a; leeDesignSpaceIntelligent2024; cardDesignSpaceInput1990). We group these design decisions into six high-level aspects of CCS that combine and interact to shape their impact: participation, inputs, curation, platform treatment, and transparency. Based on our framework, we discuss key normative implications of these systems in order to make sense of the ethical stakes involved in their design. As systems with the power to shape the information ecosystem, we emphasize three interconnected concerns: user informedness, distribution of power, and fairness. We situate CCS within the broader conversation about content moderation and establish a foundation for future human-centered research on Crowdsourced Context Systems.
2. Background
In January 2021, Twitter (now X) launched a U.S. pilot for the first large-scale CCS, which they called Birdwatch, presenting it as “a community-based approach to misinformation” (IntroducingBirdwatchCommunitybaseda). This system allowed a subset of users to suggest notes to “provide informative context” about Tweets (posts) on the platform. The same subset of users then rated these suggested notes and these ratings were used as the inputs to a note curation algorithm that aimed to select notes that would be helpful to a broad audience. In this pilot, these notes were only visible on a separate site, but Twitter announced their plan “to make notes visible directly on Tweets for the global Twitter audience, when there is consensus from a broad and diverse set of contributors.”
In March 2022, Twitter expanded the pilot, announcing that now “a small (and randomized) group of people on Twitter in the US will see Birdwatch notes directly on some Tweets” (BuildingBetterBirdwatch). Twitter was purchased by Elon Musk in October 2022, renamed X, and soon Birdwatch was renamed Community Notes. In December 2022, the pilot phase ended and users all over the world began seeing selected notes directly alongside Tweets (BeginningTodayCommunity2022). Since then, the system has continued to evolve. A key feature of X’s Community Notes is its transparency: the note curation code is open source [^1] https://github.com/twitter/communitynotes/tree/main, as are the crowdsourced notes and ratings data collected by the system [^2] https://x.com/i/communitynotes/download-data.
In June 2024, YouTube announced a pilot for a similar system, which they called Notes (TestingNewWays). In January 2025, Meta announced a pilot for their own implementation, also called Community Notes, which would launch on Facebook, Instagram, and Threads (heatheraMoreSpeechFewer2025). Finally, TikTok announced their own pilot in April 2025, which they called Footnotes (TestingNewFeature). According to the platforms, all of these new systems are based on X’s open source implementation of the note curation algorithm. Unlike X’s implementation, though, none of the new systems are open source or open data. In fact, while all of these systems serve similar purposes, they differ from each other in crucial ways due to unique design decisions made by each platform. Since these systems are increasingly prominent features of some of the most important social media platforms, it is critical to define, understand, and reason about them so the HCI community can better inform their implementation and evaluate their impact.
3. Related Work
While CCS are a relatively new innovation, these systems relate to and touch on multiple bodies of work from the HCI literature, including: research on content moderation in public discourse systems, research on fact-checking infrastructures, and research on crowdsourced responses to misinformation. We describe these areas in general terms here; later in the paper (Section 6) we provide a detailed review of the literature that is examining specific implementations of CCS.
In a democracy that values free speech, public participation in civil discussion is considered a social good (smallPolisEscalarDeliberacion2021). One classic concern of HCI research has thus been how to design systems that can support public discussion among users with diverse, possibly conflicting, perspectives (kripleanSupportingReflectivePublic2012; wohnHowHandleOnline2017). CCS are an example of such a system. Like past public discussion systems, CCS provide a form of “content moderation,” which can be used to encourage certain types of participation and discourage others (grimmelmann2015virtues). Content moderation can take different approaches: it can be “top-down,” where platforms make moderation decisions (gillespieCustodiansInternetPlatforms2018; gillespieNotRecommendReduction2022), or “bottom-up,” where users do (seeringReconsideringSelfModerationRole2020; lloyd25; lloydAIRulesCharacterizing2025; fieslerRedditRulesCharacterizing2018). Additionally, moderation can be human-powered, or use varying degrees of automation to keep up with large volumes of content (vaccaroContestabilityContentModeration2021; gorwa20). Community Notes, X’s CCS implementation, incorporates all of these approaches: it uses a centralized algorithm to aggregate moderation feedback that is sourced from users. Community Notes is an example of a bridging system, or an automated system designed to, “increase mutual understanding and trust across divides, creating space for productive conflict, deliberation, or cooperation” (ovadyaBridgingSystemsOpen2023). Indeed, according to its creators, Community Notes was designed to identify crowdsourced content (notes) “that both inform understanding (decrease propensity to agree with a potentially misleading claim) and are seen as helpful by a diverse population of users” (wojcikBirdwatchCrowdWisdom2022b). CCS thus reflect a long-standing HCI concern with designing systems that support civil discussion through content moderation, adapting these ideas to meet the scale and speed of major social media platforms.
One moderation approach frequently studied in HCI research is the use of top-down fact-checking to deal with misinformation (walterFactCheckingMetaAnalysisWhat2020; youngFactCheckingEffectivenessFunction2018). Emerging from journalistic practice, top-down fact-checking approaches empower certain individuals or organizations to verify public claims and publish determinations about their accuracy. While these techniques can have positive effects, they are time-consuming and traditionally require a large amount of professional labor (hassanDetectingCheckworthyFactual2015; junejaHumanTechnologicalInfrastructures2022), and thus may struggle to keep up with the speed at which content is published on the internet (micallefTrueFalseStudying2022). As a result, researchers have explored automated ways to scale the fact-checking process (warrenShowMeWork2025; guoSurveyAutomatedFactChecking2022; hassanQuestAutomateFactChecking). Beyond research into fact-check production, studies have also investigated the effects that fact-checks have on users who encounter them (eckerPsychologicalDriversMisinformation2022). This work has explored both the most effective ways to present fact-checks in order to shift user belief (claytonRealSolutionsFake2020), as well as the psychological characteristics that make some users reluctant to engage with them (tanakaWhoDoesNot2023). In addition to issues of efficacy, and the large cost of running such fact-check programs, social media platforms have struggled with allegations and perceptions that their fact-checks are biased against certain political ideologies (cotterFactCheckingCrisisCOVID192022; jiangReasoningPoliticalBias2020).
CCS can be seen as an evolution in content moderation that aims to address the limitations of top-down fact-checking by empowering platform users. In doing so, these new systems build on a long history of HCI research that has explored ways of harnessing the distributed wisdom of a crowd to produce accurate information in networked settings (pennycookFightingMisinformationSocial2019; heSurveyRoleCrowds2025). This research has shown that crowds can perform various tasks, such as sourcing on-the-ground information (daileyJournalistsCrowdsourcerersResponding2014; agapieCrowdsourcingFieldCase2015), or fact-checking political claims (kripleanIntegratingOndemandFactchecking2014). Studies have also found that despite being cheaper and faster than professional fact-checking, crowd-based fact-checking can be similarly accurate and effective in certain circumstances  (allenScalingFactcheckingUsing2021; martelCrowdsCanEffectively2024a; zhaoInsightsComparativeStudy2023). At the same time, crowd-based fact-checking is not a silver bullet, and other studies have shown that the composition of the crowd can affect performance (godelModeratingMobEvaluating2021), and that users may respond negatively to “social correction” that comes from the crowd (mosleh21). Crowds that are representative of the general population may be biased and cannot accurately perform every task (drawsEffectsCrowdWorker2022). Researchers have studied techniques to mitigate these biases (spiekerDiversePerspectivesCan2023) and have explored other ways that crowds can play a useful role in moderating misleading information, such as via digital juries (fanDigitalJuriesCivicsOriented2020; zhaoEffectsNonEffectsSocial2025), or via distributed moderation with up/down votes (lampeSlashdotBurnDistributed2004; epsteinWillCrowdGame2020). CCS share much in common with previously explored crowdsourced fact-checking systems, but they have a novel set of processes and structures that lead us to define them as a new class of system. Informed by these past bodies of research, we next define what makes these new systems unique.
Figure 2. Model of a Crowdsourced Context System. Users create posts, context contributors create notes about those posts, and ratings contributors create ratings for those notes. Based on these ratings, the system selects “helpful” notes, which are displayed alongside the post for users to see.:
4. Generalized Model of Crowdsourced Context Systems for Social Media
We coin the term Crowdsourced Context System to describe these social media systems, a novel and important addition to the content moderation toolbox. Such systems have several core characteristics. First, CCS rely on contributions crowdsourced from users of the platform. Second, these systems are designed to specifically elicit context from the crowd that is meant to help users interpret the original post (as opposed to other types of responses—for example, the most humorous reply). Crowdsourced context can facilitate interpretation by providing additional information that helps readers assess and interpret the veracity, potential motivations, or broader social or political context of the original post. Third, these systems supplement user content by displaying additional text, or counterspeech (pengRescuingCounterspeechBridgingBased2024a), alongside the original posts. Finally, these systems are natively integrated into social media platforms. Unlike organic crowd behavior that emerges among users, these systems are built and operated by the platforms, giving platform designers the power to influence how the systems function.
Figure 2 shows a generalized model of a CCS. Such systems are composed of several core entities, which we denote with italics through the rest of the paper. Users create posts on the platform. Contributors then provide context, in the form of text or media snippets called notes, which are associated with specific posts. At first these notes are only visible to other contributors, who assess them and produce ratings of their helpfulness. The platform uses these ratings to identify “helpful” notes, which are then published and displayed alongside the associated post everywhere that it appears on the platform, as seen in Figure 1. In implementing this general model, platforms make many individual design decisions, which we will address later.
5. Methods
To better understand CCS and their design possibilities, we performed two distinct reviews. First, we performed a systematic literature review of existing research on the first large-scale CCS implementation: Twitter’s Community Notes. Second, we performed inductive thematic analysis and affinity diagramming on a set of CCS design features extracted from documents describing the real-world CCS implementations of several major social media platforms.
5.1. Systematic Literature Review
Figure 3. PRISMA flow diagram (pagePRISMA2020Statement2021) for our systematic literature review, showing both the database search and the backwards reference search. We analyzed a final set of papers that combined the results of these two searches (n=56).:
In order to synthesize the research community’s understanding of Crowdsourced Context Systems, we performed a PRISMA-style systematic literature review (pagePRISMA2020Statement2021). We sourced relevant papers using both a database keyword search and a backwards reference search (jalaliSystematicLiteratureStudies2012). Figure 3 shows a flow diagram summarizing our search and screening process. Our goal was to identify original research that directly studies CCS implementations. Since this category of system has not been consistently named prior to our work, we focused our review on studies of the first large-scale implementation of a CCS by a major social media platform: Twitter’s Community Notes/Birdwatch. We thus used “Birdwatch” and “Community Notes” as keywords and restricted our search to articles that have been written since 2021, when Twitter first launched their implementation. Note that since this is an emerging technology and the set of articles is limited, we decided to include sources of knowledge that are valued in the HCI community but are sometimes excluded from literature reviews, such as pre-prints and non-full-length conference submissions. We thus settled on the following inclusion and exclusion criteria:
Inclusion criteria: the study is about a Twitter-style CCS, whether “in the wild” or in a controlled setting.
Exclusion criteria: The article was not written in English, does not present original research, is only an abstract/a summary of other research, or is not about the functioning of a Twitter-style CCS.
Search Strategy
We conducted our database search in November 2025. We searched seven databases that are well-known in the fields of HCI and social media research. We began with six that were used in another recent HCI literature review of social media research (zhangWhatWeMean2023): ACM Digital Library, IEEE Xplore, SAGE Journals, ScienceDirect (Elsevier), Taylor & Francis Online (Tandfonline), Emerald Insight. We additionally searched arXiv in order to also include preprints, which are common in these fields.
We searched the full-text, title, abstract, and metadata of articles using the following case-insensitive query string, and limit our results to articles published in 2021 or later:
"community notes" OR "birdwatch"
This search resulted in n=590 results.
Title Screening
We screened the search results by first considering their title. One researcher read the title for each result and removed articles that were either duplicates or were obviously off-topic, such as those that were not about social media systems. This resulted in n=254 articles.
Abstract Screening
One researcher read the abstracts of the remaining articles to further screen the results. We removed redundant articles that only summarized research that was contained elsewhere in our sample, such as books and book chapters, editorials, abstracts, keynotes, and complete conference proceedings. We also removed articles that were topically irrelevant, such as those that did not focus on a Twitter-style CCS, or that used data from such a system, but only to answer research questions unrelated to the system. This resulted in n=47 articles.
Full-Text Screening
Two researchers then read the full-text of the remaining 47 articles and applied the inclusion and exclusion criteria to screen more thoroughly. The researchers met during the process to compare their results and discuss any differences or challenging cases until they were in agreement. This produced a set of n=40 articles.
Backwards Reference Search
The first author then used this set of 40 articles as the starting set for a backwards reference search and extracted all of the n=2,207 references included in these articles. We de-duplicated these references by paper title, which resulted in a set of n=1,385 unique papers. We then performed the same screening steps that we used in the forward search: one researcher first screened the titles of the cited papers, which resulted in n=175 papers, then screened this set by abstract, which resulted in n=23, then finally two researchers performed a full-text screen. At the end of this screening process we had identified n=16 relevant papers that were not found in our forward search. We combined the sets from the two searches and produced a final set of n=56 papers.
Data Extraction
Two researchers then agreed upon data to extract from each article in the final set, which included the data source, the research questions, the main findings, and any design implications mentioned. The two researchers independently read the final set of articles and extracted this information. They then compared values, discussed any areas of disagreement, and came to a consensus for the extracted values for each paper.
Next the two researchers independently grouped similar papers into affinity groups (beyer99) based on the primary focus of the study. The researchers reconvened, compared their groups, and discussed the reasons for their groupings. Several papers were shifted between groups based on the discussion. Once the researchers agreed upon the groups, they came up with labels to describe each group based on the focus of the studies (see Table 1).
5.2. Design Feature Identification
In order to identify the design decisions that shape CCS, two authors analyzed four real-world implementations of CCS: Twitter/X’s Community Notes [^3] https://x.com/i/communitynotes/, Meta’s Community Notes [^4] https://transparency.meta.com/features/community-notes (which covers Facebook, Instagram, and Threads), TikTok’s Footnotes [^5] https://support.TikTok.com/en/using-TikTok/exploring-videos/footnotes, and YouTube’s Notes [^6] https://support.google.com/YouTube/answer/14925346?hl=en. Our analysis is based on public, company-written documentation and communications about the systems and their features, which we will collectively refer to as articles. In addition, we had access to some of the systems, which we used to complement our understanding. Unfortunately, while we attempted to become Contributors for all of the platforms, some (TikTok, YouTube) did not grant us access.
The researchers independently read through the articles one at a time and extracted design details into a spreadsheet, which resulted in a set of design features for each platform. After independently reviewing the articles for a platform, the researchers convened and performed a round of affinity diagramming (beyer99), in which they wrote the design features on sticky notes and placed them on a whiteboard. They placed similar design features together, and through this inductive method identified feature groups. The researchers then performed thematic analysis (braunUsingThematicAnalysis2006) to identify the themes of the groupings and came up with provisional group names. They repeated this process with the articles for each platform, reconvening to do another round of affinity diagramming after each round of feature extraction. These design features and feature groups were further synthesized into the aspects, dimensions, and options of the design space, which we present below in Section 7.
6. What Does the Research Community Know About Twitter-Style CCS?
Study Focus Primary Method Papers
System Impact (n=24) Data Analysis: X Community Notes Data (bouchaudAlgorithmicResolutionCrowdsourced2025b; renaultCollaborativelyAddingContext2024a; chuai25; truong2025communitynotesvulnerablerater; slaughter25; chujyo25; chuaiCommunitybasedFactcheckingReduces2024a; rollout21; kimDifferentialImpactIndividual2025e; diffusion21; chuai2025requestnoterequestfunction; matamoros25; razuvayevskaya2025timelinessconsensuscompositioncrowd; kangur2024checkscheckersexploringsource; DeepDiveXs; gao24; jama24; bobek2025communityfactchecksbreakfollower; borwankarDemocratizationMisinformationMonitoring2022a; Wirtschafter_Majumder_2023; wangEfficiencyCommunityBasedContent2024b; colemanLimitingFactorsEffectiveness2023; dekeulenaarTwitterDemotionCommunity2025; cham)
Contributor Behavior (n=16) Data Analysis: X Community Notes Data (allen22; borenstein2025; pilarskiCommunityNotesVs2023; pröllochs2021communitybasedfactcheckingtwittersbirdwatch; cikm22; kuuse25; jones22; solovevReferencesUnbiasedSources2025a; arjmandilari2025threatssustainabilitycommunitynotes; toyoda25; Phillips2025; renault25; fraxanetUnpackingPolarizationAntagonism2024c; martelPoliticalMotivesHelp2025; mohammadi2025birdwatchcommunitynotestwitter; yoon25)
System Design (n=10) User studies, design specs, algorithm analysis (mohammadiAIFeedbackEnhances2025a; wu2025crowdllmaugmentedcommunitynotes; wojcikBirdwatchCrowdWisdom2022b; hawkeye; franzmeyer2024hellofreshllmevaluationsstreams; liScalingHumanJudgment2025a; supernotes; zhang2025commenotessynthesizingorganiccomments; singh2025on; xing2025communitynotesdatasetexploringhelpfulness)
Users’ Perceptions (n=6) User studies (sharevski2025helps; liu23; shusas24; drolsbachCommunityNotesIncrease2024a; Kankham18062025; Hameleers18052024)
Table 1. The n=56 articles found in our literature review, grouped by study focus.:
Our systematic literature review of research into Twitter-style CCS found n=56 papers and identified four high-level categories of study: studies of the system’s impact, contributor behavior, alternative system designs, and users’ perceptions of crowdsourced context. These categories naturally overlap and are not mutually exclusive, but we highlight them in order to emphasize the different types of knowledge emerging from this body of research. While some papers make multiple contributions and thus span categories, we present them in Table 1 grouped according to their most significant contribution for simplicity.
The most common type of study explored the impact of a CCS on its implementing platform (n=24 papers). These studies all performed analysis of data from X’s Community Notes. These studies asked both causal questions, such as about the effect that displaying a note has on post diffusion (diffusion21; rollout21), as well as descriptive questions about the performance of X’s system in several specific use cases (chujyo25; bouchaudAlgorithmicResolutionCrowdsourced2025b). The studies found strong evidence that CCS can both reduce the spread (diffusion21) and believability (wojcikBirdwatchCrowdWisdom2022b) of misleading posts. The research also argued that the bridging-based algorithm in X’s Community Notes system has demonstrated it can identify high-quality crowdsourced notes that are likely to be perceived as helpful by users across an ideological spectrum (bouchaudAlgorithmicResolutionCrowdsourced2025b). However, the studies also showed that this bridging approach takes a long time to identify “helpful” notes for misleading posts. As a result, there is often a significant delay before notes are displayed to users (renaultCollaborativelyAddingContext2024a; chuaiCommunitybasedFactcheckingReduces2024a; rollout21). Additionally, studies have shown that the vast majority of suggested notes ($\sim 90\%$) are never published due to either a lack of ratings or consensus  (chujyo25; Wirtschafter_Majumder_2023).
Another common type of research study (n=16) examined CCS contributor behavior. While one of these studies used interviews as a data source (yoon25), all of the others relied on the open data from X’s Community Notes. The studies asked descriptive questions about how contributors to the system behaved, such as what types of posts (pilarskiCommunityNotesVs2023; cikm22) and accounts (allen22) contributors tend to leave notes on, what type of content contributors tend to include in notes (jones22; kuuse25), and what types of notes contributors rate as most helpful (solovevReferencesUnbiasedSources2025a; Phillips2025). Three key findings from these studies are that 1) the crowd exerts political bias in the posts that they select to write notes for and the ratings that they assign to notes (allen22); 2) the crowd behaves differently than professional fact-checkers (pilarskiCommunityNotesVs2023), yet often links to them as sources (borenstein2025); and 3) certain types of notes are perceived as more helpful than others, such as those that support claims with links to unbiased (i.e., assessed as politically neutral by a third-party organization [^7] https://mediabiasfactcheck.com/) external sources (solovevReferencesUnbiasedSources2025a).
A third group were design studies that either proposed or built a variation of a CCS (n=10). These papers focused on ways to improve the quantity (liScalingHumanJudgment2025a) and quality (mohammadiAIFeedbackEnhances2025a) of notes, as well as techniques to improve the performance of the note curation algorithm (hawkeye). The bridging-based algorithm for selecting “helpful” notes was introduced in one of these papers (wojcikBirdwatchCrowdWisdom2022b), as a solution to problems of bias (allen22) in a majority-rule note selection system. Recent papers have proposed ways of incorporating AI into the note-writing process and found that LLMs have the potential to generate helpful notes (supernotes; mohammadiAIFeedbackEnhances2025a).
A final group of studies sought to understand user’s perceptions of the notes produced by CCS (n=6). These studies used experiments (liu23), interviews (shusas24), and analysis of X’s data (Hameleers18052024) to answer descriptive questions about how individuals perceived the credibility and helpfulness of context provided by CCS. They found evidence that users perceive notes from CCS as just as, or even more, credible than those from professional fact checkers (liu23; shusas24). While these studies begin to explore the impacts of X’s design decisions, much is still unexplored. As we show next, the potential design space is much wider than what is covered in existing research.
7. A Design Space for CCS
Our analysis of the various CCS implementations revealed areas of common characteristics shared between the systems as well as points for variation. Analyzing these features allows us to construct a design space of possibilities. Within this design space, we identify six key aspects of CCS:
  1. (1)
    Participation: Who uses the system and what are different users are allowed to do?
  2. (2)
    Inputs: What shape do contributions to the system take?
  3. (3)
    Curation: How are suggested notes selected to be displayed?
  4. (4)
    Presentation: How are selected notes displayed?
  5. (5)
    Platform Treatment: How does the platform treat posts that have displayed notes?
  6. (6)
    Transparency: How can the public observe and evaluate the functioning of the system?
Within each aspect, we identify several dimensions (i.e., points of potential variation), which we frame as design questions, and options (i.e., potential answers for each dimension) based on the features that we observed in our analysis. In the tables in the sections that follow, we indicate when we found evidence that a platform implemented a certain option. However, platforms did not always document all of their design choices, so our tables may be incomplete. In constructing the design space, we additionally include several speculative options that we did not see in any current implementations. The sections that follow detail the dimensions and options for each aspect.
7.1. Participation
Dimension TikTok Meta Twitter/X YouTube
Who sees published notes?
All users.
Only users from a specific geo/market.
Only users that have opted in.
Who are the context contributors?
Users.
Platform employees/contractors.
Third parties.
AI systems.
Who are the rating contributors?
Users.
Platform employees/contractors.
Third parties.
AI Systems.
What must a user do to become a context contributor?
Nothing: all users are contributors by default.
Nothing: users cannot be context contributors.
Apply.
Receive an invite from the platform.
Be admitted by the platform.
Meet a quality threshold of past performance.
What must a user do to become a rating contributor?
Nothing: all users are contributors by default.
Nothing: users cannot be rating contributors.
Apply.
Receive an invite from the platform.
Be admitted by the platform.
Meet a quality threshold of past performance.
What types of eligibility requirements limit who can be a contributor?
User-age requirements.
Account-age requirements.
Lack of policy-violations.
Geographic requirements.
Language requirements.
Account authentication requirements.
Proof of some real-world credentials.
Is the number of contributors limited?
Yes: platform has discretion over who they admit.
No: all who meet eligibility criteria are automatically admitted.
How are contributors incentivized?
Paid by the platform.
Pseudonymous reputation.
Account reputation.
Unclear: contributor behavior is not visible.
How are notes and ratings solicited?
Contributors add notes to posts they organically see on the platform.
Contributors have feeds of posts that need notes, and notes that need ratings.
Platform notifies contributors of specific posts and notes to review.
Users can request notes on posts or additional review on published notes.
Table 2. Participation dimensions and options, as well as platforms that have made this design choice.:
We define participation as the design choices that shape who uses the system and what different users are allowed to do. Table 2 shows all of the participation dimensions and options. Choices that platforms make in this aspect range from those that encourage user participation in the system, to those that discourage it.
The most basic dimension is, “Who sees published Notes?” The options that platforms choose as answers to this question will greatly affect the system’s impact. The most participatory implementations like X’s allows all platform users to see these notes, while less participatory implementations limit the audience in some way beyond a user’s control, for example by their location. An even less participatory option would be to make notes an optional feature that users must opt in to see, which would greatly limit the system’s impact.
Other dimensions of participation determine who can contribute to the system, such as: “Who are the context contributors?” These dimensions shape the quality, quantity, and variety of the notes that the system produces. All implementations we examined rely on platform users as contributors, but we see some variation in what these users are allowed to contribute. For example, YouTube currently does not allow users to provide ratings; instead, those are provided by third-parties who are paid by the platform. X allows AI systems to contribute notes via an API. All of the implementations impose eligibility restrictions on users, ensuring that platforms maintain discretion over who is admitted as a contributor. Yet we can imagine a system that would automatically admit all eligible applicants, or even one with no eligibility restrictions, where all users would be contributors by default.
Besides decisions that limit the set of possible contributors, it is also important to ask questions about how platforms encourage these contributors to participate, such as: “How are contributors incentivized?” While it would be possible for platforms to pay contributors, all of the implementations that we examined rely on volunteer labor to some extent. These approaches may choose to incentivize contributions through reputation systems, for example by publicly associating contributors with either their user accounts or with pseudonyms. There is also the important question of: “How are notes and ratings solicited?” All platforms allowed contributors to add notes to posts that they organically come across while using the platform. However, this design choice could result in skewed post selection if, for example, contributors follow accounts from one end of the ideological spectrum more than the other. Some platforms may also take more active approaches that solicit notes for specific kinds of posts, or ratings for specific suggested notes, either through in-app feeds or notifications, as shown in Table 2. These approaches could be used to counter bias in the system, or to encourage notes on posts that are most likely to do harm, such as those that are going viral.
7.2. Inputs
Dimension TikTok Meta Twitter/X YouTube
What tools are note contributors given?
Free-text inputs
Ability to add source links.
Ability to add references to a portion of the post (e.g. timestamp in the video being noted).
AI note writing assistance.
API for AI note creation.
What tools are rating contributors given?
A fixed set of options (i.e., helpful, somewhat helpful, not helpful) to select from.
A free-text explanation field.
A fixed set of explanation options.
API for AI rating creation.
Which posts can have notes left on them?
All.
Posts by specific account types (public, verified, pages).
Posts of specific types of content (no ads, none that may feature minors).
Posts that reach a certain level of engagement.
Posts that have been flagged by users as needing additional context.
What restrictions apply to notes?
Must adhere to community guidelines.
Length restrictions.
Cannot be edited once submitted.
Cannot be deleted once submitted.
Source URL required.
What restrictions apply to ratings?
Must adhere to community guidelines.
Length restrictions.
Cannot be edited once submitted.
Cannot be deleted once submitted.
Can the published note on a post change over time?
Published notes can be hidden by the crowd indefinitely.
Note authors can edit or delete their notes.
Post authors can edit/remove notes on their posts.
Published notes are ”locked” after a certain amount of time.
Table 3. Inputs dimensions and options, as well as platforms that have made this design choice.:
We define inputs as the design choices that shape the type and content of contributions to the system. Table 3 shows the associated dimensions and options for the inputs dimension.
Contributions are shaped by the affordances that platforms provide to contributors to create their input, including the user interfaces for producing notes and ratings. These interfaces may allow free-form, or structured, input. Such decisions may have an effect on what information contributions contain and what they look like. For example, Meta’s implementations have a required “source URL” field, which based on the findings mentioned in Section 6, may lead to more helpful notes. The content of contributions is also shaped by additional restrictions placed on them by platforms. For example, all of the implementations that we examined required contributions to adhere to the platform’s content policy. Finally, design choices that require contributors select from a set of fixed options, may be both easier to use and produce more consistent inputs. In these ways, design decisions in this aspect will affect the volume and quality of both notes and ratings.
A final interesting question is, “Can the published note on a post change over time?”. Some implementations will give contributors the ability to edit or delete their notes, while others will prevent this. Additionally, the published note on a post can change if the curation process that selects notes (discussed below) is re-run over time. To deal with this, X’s implementation added a feature that “locks” the published note after a certain amount of time, preventing it from changing.
7.3. Curation
Dimension TikTok Meta Twitter/X YouTube
What determines which notes are shown alongside posts?
Algorithmic decision based on data from contributors.
Editorial decision made by the platform.
Editorial decision made by contributors.
What (alleged) control does the platform have over what is displayed?
Platform chooses what is or is not published.
Platform can edit note content.
Platform can remove notes if they violate platform content policy.
Platform has no control.
What data is used to make an algorithmic decision?
Ratings.
Note content (length, inclusion of a link, etc).
A note author’s past platform behavior as a contributor.
A note author’s past platform behavior as a regular User.
Demographic/personal information about the note author.
Rating contributor’s past platform behavior as a contributor.
Rating contributor’s past platform behavior as a regular user.
Demographic/personal information about the rating contributor.
What is the aim of the algorithmic decision?
Identify notes that appeals across the ideological spectrum (“bridging”).
Identify notes with majority appeal.
Identify factually-accurate notes.
Identify notes that reference reliable sources.
Table 4. Curation dimensions and options, as well as platforms that have made this design choice.:
We define curation as the design choices that affect how notes are selected to be displayed. Table 4 shows all of the associated dimensions and options. Design decisions here affect the quality of the notes that are displayed, the potential for bias in the note selection, as well as the speed and volume at which they appear.
Perhaps the most important design question is “What process determines which notes are shown alongside posts?” The existing systems that have recently grown in popularity all involve algorithmic decisions by bridging-based algorithms that, broadly speaking, identify notes that have been rated as “helpful” by contributors who have disagreed in their past ratings. Even though these bridging algorithms are currently popular, alternatives to this scheme are possible: for example, Twitter’s first implementation of Birdwatch used a majority voting algorithm. Alternatively, it could be possible to achieve bridging goals using contributor information other than their past ratings. Algorithmic approaches are also not the only option, as one could imagine a CCS that selects notes via editorial decisions by either the platform or contributors. Another possible approach is to select notes via digital juries (fanDigitalJuriesCivicsOriented2020) of users, which could be selected either randomly or based on other criteria.
A related question is, “What (alleged) control does the platform have over what is displayed?” Platforms can exert varying levels of control over the output of these systems, although current implementations all intentionally minimize the platform’s role. Still, all platforms do retain control to make editorial decisions in the case of content policy violations.
7.4. Presentation
Dimension TikTok Meta Twitter/X YouTube
Where are published notes presented to users?
Below the text of the original post.
Above the text of the original post.
In an overlay on top of the original post.
On both the original post and any re-shares of it.
Who is notified about notes when they are published?
Original post authors.
Contributors who have written/rated the published note.
Every user that replied, liked, or reposted the original post.
Every user that was exposed to the original post.
How many notes are shown on a post?
At most one.
Multiple.
How is the author of a note shown on a post?
The author of the note is anonymous/omitted completely.
The actual identity of the note author is shown.
How are a note’s ratings shown on a post?
Only whether or not it is “helpful”.
Some indicator of how helpful it is.
Table 5. Presentation dimensions and options, as well as platforms that have made this design choice.:
We define presentation as the design choices about how published notes are displayed to users. Table 5 shows all of the associated dimensions and options. Design decisions here affect the impact of a published note on users who see the associated post.
The biggest question is, “Where are published notes presented to users?” Answers to this question affect whether or not users will ever see published notes. Note presentation can range from aggressive to subtle: notes can be mandatory overlays that users must dismiss to see the associated post; notes can be shown alongside posts where they can be easily ignored; or platforms may simply indicate that a post has a note and require users to click to view them. Additionally, there is the question of which instances of the post have the published note displayed: just the original, or also re-shares of the post that appear across the platform. This will affect the number of users who have the potential of being informed by a published note. Platforms must also decide: “Who is notified about notes when they are published?” Notifications here can be used to retroactively influence the understanding of users who saw a post before it had a published note.
Other important questions affect the amount of information about notes that is conveyed to users. Twitter’s original implementation of Birdwatch showed several “helpful” notes at a time, but all of the current implementations that we examined only show one. Relatedly, platforms may provide details about a note’s author or ratings, which could inform how a user interprets a note.
7.5. Platform Treatment
Dimension TikTok Meta Twitter/X YouTube
How are posts with notes on them treated by the platform?
No different from any other posts.
Limit monetization options.
Limit positioning next to ads.
Limit spread (disabling reshares, deranking in feeds)
Increase spread (boost ranking in feeds)
How do notes interact with other content moderation features?
Replaces third-party fact checking program.
Complements platform-provided content moderation.
Notes are presented on posts alongside other “information reliability” indicators.
Table 6. Platform Treatment dimensions and options, as well as platforms that have made this design choice.:
We define Platform Treatment as the design choices about how the platform treats posts that have published notes. Table 6 shows all of the associated dimensions and options. Design decisions here affect the impact of CCS systems on the spread of posts.
The central question is, “How are posts with notes on them treated by the platform?” All current implementations make a point of emphasizing that they do not algorithmically penalize posts with notes on them (for example, by down-ranking in recommendation algorithms). But this need not be the case: perhaps certain types of notes provide a useful signal about post quality and platforms may want to suppress them. Alternatively, the platform may impose restrictions on posts with notes, like prohibiting them from being monetized or from displaying next to ads.
It is also important to ask “How do notes interact with other content moderation features?” Some platforms have shut down their fact-checking programs and replaced them with CCS. But, all of the platforms that we surveyed still have content moderation systems of some kind, which a CCS complements. Finally, YouTube and TikTok display multiple “information reliability” indicators alongside posts in tandem with notes, such as platform-provided banners or labels.
7.6. Transparency
Dimension TikTok Meta Twitter/X YouTube
How public is the note curation system?
Open source.
Open data.
Fully verifiable.
Not public.
How public is the contributor admission process?
Open source.
Open data.
Fully verifiable.
Not public.
How public is the data needed to evaluate the impact of notes on engagement?
Open data.
Available to researchers.
Available to paying customers.
Not public.
What sort of documentation about the system is public?
Blog posts.
Help articles/FAQs.
Technical documentation.
None.
Table 7. Transparency dimensions and options, as well as platforms that have made this design choice.:
We define transparency as the design choices that influence how the public can observe and evaluate the functioning of the system. Table 7 shows all of the associated dimensions and options. Design decisions here may affect the perception of the CCS by users, the public, and the research community.
Of central interest is, “How public is the note curation system?” The system’s source code and data can be public in order to allow outside study. In addition to being public, platforms could take additional steps to make their source code and data verifiable, or to allow outsiders to verify that the public code and data is indeed the same code and data that is powering the platform’s implementation. Or, platforms could do none of this and instead keep everything private, which is the approach chosen by all platforms except for X.
To fully evaluate the efficacy of such a system, we also need to ask, “How public is the data needed to evaluate the impact of notes on engagement?” This data would allow the public to holistically evaluate the CCS as opposed to other systems of content moderation. Currently this data is very hard to access, which makes holistically assessing the efficacy of these systems a challenge.
8. Discussion
Through our systematic literature review and analysis of real-world CCS implementations, we have produced a theoretical framework that can help the HCI research community better reason about, and design, these systems. In this section, we discuss future directions for CCS research in HCI. Specifically, we focus on how CCS fit within the broader field of content moderation, highlight underexplored portions of the design space, and reflect on the normative implications of CCS design choices.
8.1. Can CCS Replace Content Moderation?
CCS were created with the stated goal of addressing the misinformation problem (IntroducingBirdwatchCommunitybaseda), but it is unclear if the platforms implementing them are serious about this aim. For example, many see Meta’s choice to replace their U.S. fact-checking program with a CCS (heatheraMoreSpeechFewer2025) as a thinly-veiled effort to curry political favor with a new presidential administration that is critical of fact-checking (nixMetaEndsFactchecking2025). In fact, from a platform’s perspective, CCS have other appealing properties: they cost less than professional fact-checkers, they shield platforms from accusations of biased policy enforcement, and they offload the difficult responsibility of distinguishing fact from fiction onto users. It may be expedient for platforms to build CCS in the current moment, but it is unclear if such systems will survive if the political climate shifts. Still, as long as CCS are in a position to exert such influence over the information ecosystem, it is important for the HCI research community to study them. Most importantly, we emphasize the need to better understand how CCS perform compared to other content moderation systems. CCS may perform well in certain cases, but if on the whole they perform worse than the systems they are replacing, for example on the question of coverage, they should be used thoughtfully.
While it has been suggested that they can, CCS cannot effectively replace a fuller spectrum of content moderation tools, and were not designed to do so (elliottElonMusksMain). Content moderation is a broad term that encompasses a variety of different goals, including preventing crimes, mitigating lawful but awful behavior, and promoting high quality content (grimmelmann2015virtues). The tools used as a part of content moderation are similarly varied and include both “hard” approaches, like content take-downs and account bans, and “soft” approaches, such as content labeling (zannettouWonElectionEmpirical2021). Systems that provide context are narrow in their their goal and mechanism: they are a soft moderation attempt to aid content interpretation by providing users with additional information (matamoros25). And yet, the empirical results summarized in this paper demonstrate that CCS may play a meaningful role in the content moderation ecosystem, and it is worth exploring how they may integrate with other parts of it. For example: how might CCS be integrated into platforms’ automated moderation systems? In the Platform Treatment design aspect, none of the implementations that we examined used the CCS output to affect the visibility of posts. Perhaps these systems can be integrated into platform content recommendations algorithms to promote or demote posts based on their notes.
8.2. Underexplored CCS Design Aspects
Our design space reveals several ways forward for a HCI research agenda on CCS. Several aspects are underexplored by current research, particularly the Participation aspect. Encouraging positive participation in collaborative computing systems is a classic topic in HCI research (krautBuildingSuccessfulOnline2012) and one in which the community could make strong contributions. One dimension of participation which has not yet been explored is how platforms may play a more active role in soliciting notes and ratings from contributors. Past HCI work on how to best make use of the various skills of crowds may be instructive here (bernsteinSoylentWordProcessor2015; kitturFutureCrowdWork2013). It is worth noting that the more consequential these systems become, the more they will be targets for negative forms of participation as well, primarily gaming (danielQualityControlCrowdsourcing2018; epsteinWillCrowdGame2020). Bridging algorithms are designed to be resistant to certain types of gaming (wojcikBirdwatchCrowdWisdom2022b), but vulnerabilities may emerge. One other area for study is the Presentation aspect, specifically how different approaches to visually displaying notes may affect user understanding of misleading claims (gamageLabelingSyntheticContent2025; martelMisinformationWarningLabels2023). Another promising area is the Inputs aspect, where recent work has begun to explore the ways in which LLMs can supplement human effort and increase the quality and quantity of notes (supernotes; mohammadiAIFeedbackEnhances2025a). Of particular interest is X’s recent decision to allow AI note contributors via a new API (liScalingHumanJudgment2025a). How such a decision affects the quantity and quality of notes on the platform should be a near-term focus of HCI research. Of course, studying how changes to CCS impact information dynamics across a platform will be difficult to study without more innovations in the Transparency aspect. X’s approach to transparency in their Community Notes system has been a boon for studying CCS. Hopefully other platforms will soon adopt similar approaches to support ongoing studies of CCS effectiveness.
8.3. Normative Guidelines for CCS Design
CCS are complex sociotechnical systems (chernsPrinciplesSociotechnicalDesign1976) that integrate people, algorithms, and speech. These systems’ growing influence over the information shown on major social media platforms means that the design choices made in their implementation can have significant social consequences (friedmanBiasComputerSystems1996; flanaganEmbodyingValuesTechnology2008). With such high stakes, it is crucial that a human-centered research agenda consider the implications of different design choices, in order to holistically evaluate these systems’ impacts. One way to do so is to identify the primary normative concerns of such a system. We emphasize three interconnected concerns that should guide CCS design: user informedness, distribution of power, and fairness.
8.3.1. User Informedness
While social media serves many different purposes, CCS are specifically concerned with social media’s ability to inform the public. Such systems should thus be evaluated by how well they facilitate user informedness: for example, the degree to which they inform users about the accuracy of information, minimize the spread of misleading posts, or otherwise improve information quality on a platform. Our literature review surfaced ample evidence that the notes produced by CCS do have the power to increase user informedness, both by decreasing an individual’s belief in misleading information, and by decreasing the spread of this information (wojcikBirdwatchCrowdWisdom2022b; renaultCollaborativelyAddingContext2024a; slaughter25; chuaiCommunitybasedFactcheckingReduces2024a; rollout21; diffusion21). Additionally, there is evidence (kimDifferentialImpactIndividual2025e) that these notes may produce less of a backfiring effect than forms of social correction that is not mediated by a CCS (mosleh21). But while published notes have been shown to be effective at promoting understanding, we must also question how effective CCS are at producing these notes. Several of these same studies found that Twitter’s implementation of CCS produces notes too slowly to prevent the spread of the most viral content (rollout21; chuaiCommunitybasedFactcheckingReduces2024a; renaultCollaborativelyAddingContext2024a). There is also a significant gap in the research on the question of coverage, or how many of the misleading posts on a platform receive “helpful” notes, which is currently unanswered. The research suggests that there is a CCS design trade-off between speed/coverage and quality/helpfulness: X’s implementation has optimized for the latter, perhaps at the expense of the former. Whether this is the right decision should be up for debate.
For designers of CCS hoping to improve user informedness by increasing the system’s speed and coverage, there are several aspects from our design space that could be relevant. For example, within the curation aspect, designers could explore bridging algorithm changes that could lead to faster identification of “helpful” notes. Additionally, design changes within the participation and inputs aspects, such as those that would allow a larger contributor base or that would actively solicit notes on the most viral posts, could lead to a higher volume of notes and ratings, and thus quicker selection of helpful notes. When it comes to the difficulty of assessing the coverage of CCS, design changes within the transparency aspect could make this easier to study, perhaps by providing easier access to post engagement metrics.
8.3.2. Distribution of Power
One of the selling points of CCS is that they shift power over content moderation from the platform to the community. Giving users more of a say over the content that they see on their platform can make moderation decisions more legitimate and trusted (shusas24; liu23; drolsbachCommunityNotesIncrease2024a), which is an admirable goal. Yet such claims require further scrutiny, and CCS should be evaluated to the extent that they actually empower users and do not merely allow platforms to shirk responsibility for misleading content. Evidence from our literature review suggests that bridging-based algorithms, by design, fail to produce notes for the most polarizing content (bouchaudAlgorithmicResolutionCrowdsourced2025b). For example, X’s implementation of the bridging algorithm requires polarized contributors to agree that a note is helpful in order to publish it. The editorial function of the bridging algorithm can thus limit the type and scope of contributions in two important ways: in terms of framing, the algorithm may prefer centrist points of view and content that could be agreeable to both sides of an issue; in terms of selection, agreement on a note may not occur for contentious topics, preventing its publication. Whether or not users at the ends of the ideological spectrum should be empowered to make moderation decisions is a valid question, but it is uncontroversial that the most polarizing content needs moderation, too. Research has found evidence of harmful content being handled poorly by the system, either because potential notes never reach helpful status (chujyo25), or because notes that do are not actually helpful for certain types of content, such as disinformation that is couched in humor (matamoros25). Zooming out, the low publishing rate for suggested notes (around  10% on X (Wirtschafter_Majumder_2023) and  6% on Meta [^8] https://x.com/guyro/status/1965800300557857053) gives us pause that the system is actually empowering these note authors. A related, underexplored area for research involves studying which note contributors are empowered by CCS. This research area is difficult to study holistically, as these systems are pseudonymous, but some work has begun to explore it qualitatively (yoon25) and via partnerships with platforms (allen22). More research is needed to understand how CCS can empower users of certain demographics or political views over others.
Our design space offers several ways that designers can further empower users via CCS. First and foremost is innovation in the participation aspect that would increase the set and diversity of contributors. The challenge for the platforms will be ensuing that quality remains as the participating contributor base increases, which is a tension that has been thoroughly studied in the crowdsourcing literature (danielQualityControlCrowdsourcing2018), for example. Innovation in the inputs aspect may increase the quality of notes across the board, perhaps by giving note authors helpful feedback (mohammadiAIFeedbackEnhances2025a), empowering more users to make successful contributions to the system. Another route to shift power to users is to disempower the platform. Though CCS are often described as ways to reduce platforms’ centrality in content moderation, platforms still exert power over the system in important ways, such as by deciding who is admitted as a contributor. Platforms could further guarantee their neutrality through changes to the transparency aspect, such as by making their contributor selection process public. But considering that X’s CCS implementation is the only even partially open system that we examined, a first step is for other platforms to follow X’s lead and open source their note curation algorithm and the crowdsourced data that they collect.
8.3.3. Fairness
CCS are often touted for their potential to implement content moderation in an unbiased way. Yet these systems are not inherently free from bias. Results from our literature review detail bias in contributor behavior, in terms of the types of accounts that they leave notes on, as well as how they rate suggested notes (allen22). Bridging algorithms have helped mitigate bias in the note curation process, with the caveats noted above, but additional bias can still be introduced—for example, in the note creation process, or in the selection of contributors. To give a concrete example: X’s bridging algorithm finds agreement among contributors, not the entire user base. If contributors’ beliefs are not representative of users’ beliefs then the algorithm’s outcomes may be a biased representation of the community, and not necessarily have broad appeal among users. There is some evidence that so far X’s implementation has not suffered from this problem, but it is worth considering. Also, accusations of bias in top-down moderation are often due to the opacity of such approaches and may be mitigated by transparency efforts. Only X’s implementation has provided any transparency, and yet the lack of transparency about the contributor base has made analyzing them for bias impossible. Finally, the approach of displaying context rather than removing posts is arguably fairer in that it lessens the damage from false positives, by empowering users to come to their own conclusions about content’s veracity (pengRescuingCounterspeechBridgingBased2024a). However, there is a trade off: leaving up questionable content, even with helpful context, can potentially do harm to those targeted by the content.
Designers should consider several aspects from our design space that can influence the fairness properties of a CCS. First off, all platforms have taken the approach that it is fairer not to remove misleading content. However designers may explore other options from the platform treatment dimension to identify certain cases where the platform should take action on content flagged by the crowd as misleading. Second, current implementations tend to display a single, definitive note, even if there may be several that are determined to be “helpful”. The presentation aspect offers alternatives that could further the aim of allowing users to come to their own conclusions, such as by showing multiple notes when there are multiple with helpful ratings. Additionally showing information about the note author could be helpful, as it could help users identify the credibility of the source. This may be at odds with contributor privacy, which current implementations have opted to protect. Finally, innovations in the curation aspect could explore other methods of selecting notes to display that may be regarded as fairer. Current implementations of the bridging algorithm use past ratings to identify contributors with diverse perspectives: perhaps it would be fairer to instead use in-app behavior? This could provide a fuller picture of a contributor’s ideological leanings. Encouraging more transparency would also give assurances of fairness, both in how contributors are selected, and how notes are published.
8.4. Limitations
Since there was no agreed upon name for CCS before our paper, we decided to focus our literature review on studies of Twitter-style CCS, and thus may have missed earlier work that would have meaningfully informed the design space. Our analysis of platforms’ implementations of CCS is based on public documentation and may not perfectly reflect the actual features of the deployed systems. Finally, our design space includes several non-exhaustive, speculative options that were not seen in any implementations, but nonetheless are included to illuminate alternative possibilities.
9. Conclusion
Crowdsourced Context Systems are increasingly prevalent in the content moderation strategies of prominent social media platforms. To understand these systems, their impacts, and how they may be better designed, we conducted a systematic literature review of existing research studying implementations of CCS and analyzed several major social media platforms’ implementations of such systems. Despite recent trends towards the adoption of CCS, we emphasize that they are only one piece of a larger content moderation suite, and suggest against using them as drop-in replacements for traditional fact-checking approaches. The framework we produced can be used to reason about these systems, and consists of a theoretical model and a design space. We additionally discuss some of the key normative concerns that should be considered when designing CCS. Such systems have potential to address the challenge of misinformation, but to realize this potential, future design research from the HCI community must focus on ways to understand and improve these systems’ performance.